A Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de mathématiques.

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.10

Title word cross-reference

(T, M, S) [267]. 1 + ϵ [386]. 2 [368, 377]. 4 [380]. 9 [392]. [−1, 1] [275]. [αn] [213]. A(K) [346]. A_{p}^{r}(G) = A_{p} ∩ L^{r}(G) [388]. A_{R}(K) [346]. B [93, 98, 313]. (U(p, q), U(r, s)) [279]. C [80, 142]. C(α) [266]. C(K) [185, 344]. C^{∗} [282]. C_{R}(K) [346]. F [155]. f^{n}(x) = f(x) [181]. F_{4} [268]. G_{2} × PU_{3}(D) [258]. B [203]. GL(3) [316]. GL(3, q) [108]. GL(4, q) [108]. G ⊆ SU_{2}(C) [304]. GU(1, s) [347]. h^{1} [324]. H^{∞} [259]. j [392]. K [349, 303]. L [300, 389, 286, 322, 334]. L^{p} [318, 275, 375]. λ(τ) [139]. μ(z, β, α) [241]. n [101, 59, 63, 211, 188, 375, 213]. O(2n) [397]. p [316, 179, 397, 367, 334, 258, 280, 357]. PGL(3, q) [108]. PGL(4, q) [108]. q [391, 89, 392]. Q_{α}(R^{n}) [370]. S [177]. S^{1} [365]. SL_{2} [377]. SO(2n) [397]. Sp_{2}(F) [357]. U(n, n) [260].

-Adic [258, 316, 397, 203, 334, 357]. -Algebras [282]. -analogue [392]. -ary [313]. -convex [80]. -dimensional [377, 59, 375]. -equation [155]. -function [375]. -functionals [349]. -Functions [300, 286, 177, 389, 322]. -Groups [280]. -Laplacian [367]. -manifolds [380, 365]. -Nets [267]. -space [101]. -tensors [179]. -Theory [303]. -ultraspherical [391]. -values [334].

3-manifolds [360].

Abelian [296, 278, 163]. absolute [125]. absolutely [129, 50]. abstract [30, 74, 41]. abstrait [41]. additive [174, 393]. Adic [258, 316, 397, 203, 334, 357]. Affine [304, 158, 90, 208, 48, 338]. AI [398]. aide [379]. Algebra [299, 246, 288, 86, 177, 399, 180]. Algebraic [273, 339, 317, 88, 4]. Algebras [272, 282, 274, 287, 280, 93, 98, 217, 135, 388, 58, 398, 377, 333, 321, 254, 354, 215, 153, 176]. Algèbres [254]. Algorithm [239, 61, 110]. Almost [242]. alternative [145]. ambiguity [250]. amenability [354]. Ampère [331, 325]. ampleness [371]. analogue [392]. Analyse [132]. Analysis [281, 312, 132]. Analytic [170, 259, 95]. Angular [21]. anneau [41]. annulus [169]. aposyndetic [323]. Application [286, 332]. Applications [265, 27, 385, 129, 185, 351, 321, 89, 257, 369]. approach [317]. Approximation [349, 330, 57, 60, 257]. Approximations [59, 382]. arbitrary [6]. arc [157]. area [59]. arithmetic [72, 42]. Arrays [267]. Arthur [379]. ary [313]. Aspects [272]. Associated [280, 189, 394, 390, 64, 16]. Association [267]. Asymptotic [262, 384, 231, 233, 234, 314, 257, 241]. Asymptotics [390]. Aubry [123]. auteurs [311]. Author [311]. Autometrization [200]. Autometrized [93, 98]. Automorphic [292, 286]. automorphisms [74]. Averages [275, 349, 92]. Axiomatic [28]. Axiomatique [84]. Axioms [140].

balanced [208]. ball [375]. Banach [253, 276, 388]. Bargmann [285]. barrier [54]. base [143]. Bases [303]. BBP [313]. be [167]. Behavior [234, 314]. behaviour [257, 241]. Bergman [335]. Bernstein [226, 183]. Bessel [383, 91, 182]. between [197, 24, 55, 153]. Bi [283]. Bi-Period [283]. Bianchi [46]. bidisk [335]. Biflatness [354]. biinfinis [248]. bilateral [224]. Blaschke [271]. block [208]. BMO [353, 382]. BMO-estimates [382]. bodies [381]. Boolean [86, 216, 93, 98, 180, 215, 188]. boundaries [161]. Boundary [189, 374, 375]. bounded [9, 12]. Boundedness [14]. bounds [360]. Branching [316]. Brascamp [355]. Brownian [281]. Brunn [270]. bundles [371].

calculus [156, 338]. can [167]. Canonical [376, 361, 151]. Carleson [353]. carpet [390]. Carpets [281, 297]. carrés [254]. Case [258, 396, 336]. Cayleyan [3]. central [334]. Centre [288]. Certain [97, 7, 124, 135, 55, 164, 112, 72, 121, 64, 208, 235]. certains [121]. chain [219]. chains [148]. changes [124]. channel [54]. Character [366]. Characterisation [332]. Characteristic [298, 32]. characterization [216, 236]. Characterizations [391, 381, 161]. Characters [256, 399, 202]. Chowla [337]. Chung [165]. class [130, 322, 52, 335]. classes [90]. Classification [335, 330, 398, 250]. closure [150]. CM [396, 322]. CM-case [396]. Coactions [282]. coefficient [169]. Coefficients [234, 139, 324]. cognate [224]. Cohomology [289, 373, 217, 328, 345]. Colmez [337]. colour [175]. colourings [332]. combinations [226]. Combinatorial [300, 44, 27]. commutative [339, 106, 254]. commutatives [254]. commutativity [350, 356]. Compact [285, 185, 95, 325]. compactes [185]. compactness [90]. comparability [236]. Competing [269]. Complete [246, 314, 398]. Completely [11]. Complex [289, 257, 325]. complexes [400]. Composite [263]. Computing [340]. Concerning [209]. condition [164]. conditionally [393]. cone [244]. configuration [116, 133]. Configurations [308, 133]. conformal [220]. Congruence [24, 211]. congruences [221, 51]. Conjecture [251, 363, 386, 396, 165, 337]. conjugate [5]. conjugates [164]. connected [119]. connection [153]. connections [362]. consecutive [71]. consequences [373]. constituents [319]. constraints [122]. constructions [378]. Contact [299]. Contactomorphisms [290]. continua [82, 323]. continued [133]. continues [133]. Continuous [181, 10, 391, 39]. Contribution [70, 70]. Contributions [136]. controlled [401]. Convergence [263]. convergent [129]. Converse [275]. Convex [290, 312, 5, 112, 90, 353, 57, 381, 80]. convexity [359]. Convexly [263]. convolution [393]. correspondence [319]. countably [107]. Counting [294]. courbure [1, 63, 149]. coverings [190]. Covers [296]. Coxeter [310, 304]. Criteria [371]. criterion [192, 150]. critical [315, 6, 369]. Csima [239]. cubic [384, 332, 351]. Curvature [360, 314, 21, 1, 196, 378, 63, 149]. curve [334]. Curves [298, 101, 221, 78, 369]. Cyclic [131, 158, 110, 100]. cyclotomic [130].

D. [177]. dans [132]. Darboux [129]. d’Arthur [379]. decimals [137]. decomposable [330]. Decomposition [287, 323]. Decompositions [247, 68]. Dedekind [261]. defined [195]. definition [325]. deformations [340]. Degenerate [367, 362]. degree [30, 154, 9]. degrees [50]. density [22]. d’entrelacement [379]. denumerable [143]. depth [54]. Designs [249, 243, 208]. d’espaces [185]. determinant [321]. determined [228]. Diagonal [243, 384, 135]. diagrams [43]. difference [200, 210, 152, 209]. Differentiable [220]. differential [128, 120, 75, 318, 184, 121, 64, 167, 16, 223]. différentiels [121]. Dilatations [271]. dilated [349]. Dimensional [277, 183, 377, 59, 8, 375]. Dimensions [307, 297, 63]. Diophantine [56, 60]. Direct [25, 109]. Dirichlet [352]. Discrete [282, 115, 397, 2, 391]. disjoint [13]. disk [352]. disks [374]. Distance [104, 93]. distance-theoretic [93]. distribution [237, 327, 225, 69, 343, 132]. distributions [97, 132]. distributive [85]. distributives [84]. Ditkin [388]. divergent [172]. division [114, 176]. divisors [358]. domains [348, 57]. doubling [382]. Dual [279, 258, 52]. duality [101]. d’un [41, 63]. duodenary [206]. dynamics [47, 83]. Dynkin [304].

edge [332]. edge-colourings [332].

eigenfunctions [31, 18]. eigenspaces [318]. eigenvalues [352, 320]. Einstein [196]. Eisenstein [260]. Elementarfunktionen [49]. Elementary [4, 42, 49]. Elements [339, 303, 141, 138]. eleven [100]. ellipsoid [59]. elliptic [315, 222, 141, 140, 334, 369]. Embeddings [277]. Endoscopic [300]. ennuple [221]. Ensembles [248]. enumeration [240]. Enumerator [262]. Enveloping [58]. Epstein [26]. equation [155, 189, 222, 181, 55, 184, 167, 16, 50, 77, 77]. Equations [295, 384, 120, 75, 318, 162, 184, 56, 105, 121, 64, 324, 223, 178]. equivalence [90]. equivalence-classes [90]. equivalences [380]. equivariant [328]. Ergodic [294]. Erratum [327]. error [124]. espace [63]. espaces [132]. essential [356]. Estimates [264, 324, 382]. Eta [261]. étude [70]. Euclid [61, 110]. Euclidean [141, 305]. Exceptional [273, 268]. Excluding [313]. Existence [309, 143, 208]. Expansions [231, 233]. explicit [351, 378, 392]. exponent [315]. exponential [88]. expression [392, 235]. Extension [190, 144]. Extensions [217, 362, 364]. Exterior [271, 286]. Extrema [352]. extremal [157]. extremals [381]. Extreme [126, 293, 244, 206]. extremely [388]. Extremum [38].

facially [65]. facially-regular [65].

Factorization [31, 348, 40]. factorized [167]. Factors [283, 160]. faiblement [185]. families [369]. family [174, 143]. Fano [376]. Farey [89]. FC [219]. FC-chain [219]. Fibers [290, 348]. field [130, 316, 199, 83, 162, 88, 37, 76, 357]. Fields [252, 17, 61, 110, 351, 377, 32, 322]. Finding [313]. Finite [159, 94, 271, 277, 7, 199, 162, 15, 54, 218, 250, 321, 203, 8, 40, 50, 176]. finite-dimensional [8]. finitely [330]. Finsler [317, 378, 63]. finslérien [63]. First [299, 279]. Flächen [49]. flag [378, 338]. Fleming [387]. flow [385, 232, 390]. Flowers [238]. Focal [218]. force [17]. force-fields [17]. Form [285, 146, 206, 83]. Formal [348]. Forms [292, 305, 126, 144, 95, 29, 163, 167, 336]. Formula [283, 394, 337]. Formulae [262, 313, 384, 27]. formulations [86]. Four [307, 34]. four-space [34]. Fourier [155, 225, 97, 139]. fractal [370]. fractional [361, 73]. fractions [133]. Fréchet [253, 12, 68]. Fréchet-Like [253]. Free [383, 362, 15, 380, 393]. French [96, 185, 1, 41, 379, 63, 149, 254, 121, 132, 70, 133, 77]. frequencies [39]. Frobenius [201]. Function [228, 194, 261, 124, 351, 139, 14, 16, 375, 370]. function-lattices [14]. functional [49, 318, 181]. Functionals [12, 186, 349, 227, 381]. Functions [300, 263, 286, 295, 49, 187, 327, 177, 225, 394, 5, 170, 55, 150, 389, 174, 390, 342, 113, 331, 95, 399, 39, 173, 322, 26, 134, 72, 353, 182, 169, 143, 257, 64, 188, 207]. Functoriality [361]. fundamental [365, 380, 93]. Funktionentheorie [49].

Galois [361, 291, 395, 106, 153]. General [224, 350, 162, 55, 325, 35]. Généralisation [63, 149]. generalization [116, 75, 26, 149, 63]. generalizations [27]. generalize [184]. Generalized [47, 8, 318, 46, 62, 196, 3, 375, 161]. Generated [301, 205]. Generating [399, 182]. generic [319]. genus [99, 191]. geometries [159, 45, 48]. geometry [128, 339, 141, 140, 127]. German [49]. given [122, 30, 74, 320]. Glaisher [198]. global [359]. Goresky [251, 338]. Gotlieb [239]. Gradings [272]. graph [154]. Graphs [30, 296, 304, 236, 332, 40, 160]. gravitation [46, 196, 178]. gravitational [105]. Green [95, 64, 16]. group [118, 98, 30, 74, 365, 217, 219, 87, 138, 322, 166, 202, 380, 13, 43, 50]. groupoids [345]. Groups [301, 282, 273, 164, 285, 310, 9, 302, 280, 339, 119, 329, 15, 218, 342, 367, 203, 204, 205, 102].

H [165, 367]. H-type [367]. Hamiltonian [47, 83]. Hankel [336]. Hardy [315, 367, 335, 336]. Harmonic [281, 271, 179, 132]. harmoniques [132]. Hausdorff [297, 142]. heat [390, 184]. Hecke [399, 377]. Heegner [364]. Heisenberg [274]. Helly [57]. help [379]. Hereditarily [276]. Hermann [26]. Hermite [255]. Hermitian [320]. Heyting [246]. higher [336]. Hilfsmittel [49]. Hill [77]. Holomorphic [374]. Holomorphie [379]. Holomorphism [379]. Homeomorphic [259]. homogeneous [373, 323]. Homological [272]. Homology [296]. homomorphic [135]. Homotopy [265, 380]. Hopf [288]. Hulls [290]. Hyperbolic [312, 360, 179]. hyperelliptic [151]. hypergeometric [224]. Hypergroups [256]. hypersurface [78].

Ideal [259, 361, 398, 37, 136]. Identities [249, 51, 46]. Identity [300, 382]. II [237, 98, 396, 62, 266, 109, 233, 79]. III [373]. immersibility [87]. implicative [216]. Incidence [79, 131]. incomplete [208]. Indecomposable [276, 395, 11]. indefinite [144]. independent [69]. indeterminates [237]. Index [311, 15, 157]. Induced [282, 287, 67]. Inequalities [275, 229, 367, 230]. inequality [370]. inertia [320]. Infeld [167]. infinite [28]. Infinitely [295]. Injectivity [246]. Integers [310, 22, 71, 213]. Integral [187, 21, 55, 62, 203, 2, 356]. integrals [10, 129]. integration [207]. interacting [387]. Interior [324]. interlacing [379]. Interpolants [293]. intersection [112]. interval [236]. Invariant [278]. Invariants [273, 308, 94, 67]. inversion [171]. involution [100]. Irrationalities [261]. irrationality [192]. irreducible [237, 330, 319, 166, 13, 50]. irregular [170]. isometric [358]. isometries [358]. Isomorphism [308]. Isoperimetric [156]. isotherm [1]. isotherme [1]. Iterated [147]. Iterates [73]. iteration [72]. Iwahori [377].

J. [165]. Jacobi [197, 75]. John [370]. Joubert [255].

Kähler [314, 325]. kinematics [141]. knots [36]. Knotted [296]. Köthe [227]. Kottwitz [251]. Kummer [364].

ladders [31]. Lagrangian [374]. Langlands [286, 319]. Laplace [189, 18, 171]. Laplace-operator [18]. Laplacian [367, 352]. Latimer [23]. Lattice [262, 214]. Lattices [74, 193, 84, 6, 85, 14, 153]. Laurent [234]. laws [383]. Lebesgue [96]. left [68]. Lehmer [396]. length [228]. levelling [228]. LF [400]. LF-spaces [400]. Lie [217, 329, 285, 342, 58, 333, 280, 375]. Lieb [355]. Like [253]. Limit [292, 393]. limits [68]. line [127]. linéaires [185]. lineal [141]. Linear [226, 301, 229, 190, 120, 75, 394, 122, 185, 115, 227, 211, 230, 48, 153]. linear-space-valued [190]. linearly [170]. Liouville [32]. Littlewood [177]. LMO [324]. Local [299, 359, 377]. Locally [395, 344, 40]. locus [347]. Lorentz [2]. low [329, 352].

MacDuffee [23]. Machin [313]. Machin-Type [313]. MacPherson [251, 338]. Manifolds [289, 326, 360, 365, 250, 18, 380, 242, 161, 325]. Many [295, 245]. Map [175]. Map-colour [175]. mapping [135]. Mappings [271, 190, 401]. Maps [306, 259, 359, 329, 52]. Mathieu [102]. matric [135]. matrices [245, 135, 131, 320, 34]. matrix [24, 201, 235]. matroid [372]. Maximal [232, 259, 382]. maximum [69]. McCoy [201]. Mean [10, 1]. Mean-continuous [10]. measurable [170]. Measure [168, 21]. measures [353, 382]. mechanical [4]. mehrerer [49]. Mendelsohn [249]. Meromorphic [257, 95]. Method [286, 229, 230, 167]. methods [195, 25, 109]. metric [244]. metrics [314, 378]. Metrization [103]. Meyer [144]. Minimal [56, 374]. Minkowski [270]. minuscule [373]. Model [269]. Modular [118, 295, 204, 139]. modules [340, 333, 11]. Moduli [368, 366]. Moments [369]. Monge [331, 325]. Monotonically [401]. Morphisms [265]. Motion [281, 35, 105, 19]. motions [98]. mots [248]. Moufang [268]. moyenne [1]. multicoherent [79]. multigroups [45]. Multiple [394, 49, 320]. multiplication [335]. Multiplicative [358, 186]. Multiplicities [292]. Multiply [173]. mutables [49]. mutation [387]. Mutually [323].

Nakajima [359]. nature [120, 389, 235]. negative [187]. Nets [267, 94]. network [232]. Neumann [352, 321]. neutral [318]. Nevanlinna [293]. Newtonian [17]. Nikodym [96]. Nilpotency [280]. Nilpotent [303]. Nirenberg [370]. Non [291, 106, 129, 400, 208, 382]. non-absolutely [129]. non-commutative [106]. non-doubling [382]. non-existence [208]. Non-normal [106]. non-semisimple [106]. non-separable [400]. Non-Surjective [291]. noncommutative [340, 136]. Noncompact [264]. nonexistence [7]. Normal [122, 148, 137, 179, 106]. normalisés [379]. normalized [379]. Normals [253]. Note [17, 137, 211, 204, 64, 66, 3, 26, 172, 117]. nouvelles [178]. Null [223]. number [130, 214, 162, 88, 42, 188]. Numbers [298]. Numerical [207, 94].

Occurrence [279]. odd [329]. ODEs [327]. Old [36]. one [154]. one-regular [154]. onto [135]. open [400, 180]. opérateurs [379]. operations [86]. operator [18, 3, 325]. operators [330, 367, 379, 335, 356, 382]. orbifold [345]. order [326, 327, 120, 101, 73, 16, 50]. Ordered [267, 115]. ordinary [224]. origin [37]. Orthogonal [243, 34, 267, 327, 150, 221]. oscillatoires [178].

Packing [297]. Pair [258]. Pairs [279, 384]. paper [66]. Pappus [116]. Parabolic [278, 324]. paracompact [107]. parallelogram [214]. parameters [379]. paramètres [379]. Partial [256, 184, 69]. partially [339]. particles [35]. partitions [199]. Paths [238]. perfect [117]. Period [283, 100]. Periodic [48, 327, 121]. périodiques [121]. perturbation [321]. perturbed [121]. perturbés [121]. Pick [293]. Plane [298, 54, 220]. planes [7, 158]. Point [262, 6]. points [364, 214, 170, 220]. Poisson [375]. Polar [247]. Poles [286, 260, 257]. polyhedra [65, 111, 38]. polyhedron [355]. polynomes [41]. Polynomial [290, 273, 197, 41, 392]. polynomials [312, 91, 226, 183, 237, 327, 391, 182]. polytope [372]. positive [314, 339, 32, 235]. possible [167]. postage [70]. postage-stamp [70]. poste [70]. Postulates [85, 215]. potential [315]. pour [311]. power [33, 89]. powers [24]. prescribed [257]. preservers [350]. presheaves [340]. prime [42, 213]. prime-number [42]. Primes [295, 329]. primitive [214]. principal [316, 319, 63, 149, 357]. principle [326]. problem [57, 169, 142, 70, 133, 4, 77]. problème [70, 133, 77]. problems [44, 189, 156]. processes [387]. product [33, 13]. Products [284, 138]. progressions [42]. Projective [289, 308, 159, 7, 48]. Projectors [318]. proof [146, 42]. Properties [247, 253, 93, 129, 112, 18, 334, 14, 38, 80]. property [398, 89, 142]. pseudo [354]. pseudo-amenability [354]. Pseudolocality [385]. pseudometrics [190]. pseudopowers [343].

Quadrangles [268]. Quadratic [310, 261, 199, 61, 144, 29, 334, 163]. Quantum [373, 333, 4]. Quartic [298]. quasi [350, 254, 254]. Quasi-commutative [254]. quasi-commutatives [254]. quasi-commutativity [350]. Quasiconformal [290]. Quasiconvex [81]. Quasigroup [249]. Quelques [149]. Quiver [265]. Quotient [276, 256]. Quotients [264].

Radiation [105]. Radical [278].

Ramanujan [51]. ramified [316]. random [69]. rank [371, 329, 342, 368, 320, 28]. rays [244]. real [61, 342]. Realizations [306]. recombination [387]. reconnaissables [248]. recurrence [394]. recursions [119]. Reducibility [346, 357]. reducible [22]. reducing [335]. Reduction [253]. Reflection [301, 302]. reflections [205]. reflexive [368]. regions [90]. Regular [191, 306, 65, 111, 99, 154, 180, 169, 38].

Related [297, 331, 182, 370]. relation [197, 55]. relations [24, 79]. relatively [213]. Relativity [35]. relaxation [229, 230]. remark [143]. remarks [149, 202]. remarques [149]. renormings [344]. Representation [227, 29, 288, 171]. Representations [291, 265, 274, 287, 316, 118, 395, 397, 9, 71, 203, 204, 166, 13, 67, 108, 50]. residue [210]. residues [89]. resolvable [208]. restrictions [397]. results [331, 332, 36]. Reversibility [387]. Revisiting [359]. Ricci [314, 362, 385]. Riemann [146, 49, 362, 195]. Riemannian [326, 179, 18, 221, 242]. Riemannschen [49]. rigidity [331]. Ring [288, 86, 328, 41]. Rings [272, 145, 216, 135, 114, 106]. Robinson [66]. Roch [146]. Root [307]. roots [245]. rotund [344]. ruled [371]. rules [316].

S. [363]. samples [122]. Satisfying [295, 55]. scalaire [63, 149]. scalar [378, 63, 149]. Schemes [267]. Schrödinger [167]. second [326, 327, 120, 16]. second-order [16]. sector [68]. Segal [285, 354]. Selberg [337]. selection [387]. Self [329, 380, 52]. self-dual [52]. self-equivalences [380]. Self-maps [329]. Semi [304, 115, 58, 39, 373]. Semi-Affine [304]. semi-continuous [39]. semi-ordered [115]. semi-simple [58]. Semi-simplicity [373]. semidefinite [235]. Semigroup [272, 87]. semigroups [168]. Semilattices [284]. semilinear [315]. semisimple [342, 106]. Separable [253, 400]. sequence [266]. sequences [394, 89]. Series [260, 224, 316, 33, 218, 397, 319, 172, 223, 357]. set [170]. Sets [246, 297, 197, 388, 81, 104, 112, 210, 6, 138, 152, 209, 180, 28, 80]. several [237, 207]. Shahidi [286]. sheaves [340, 368]. Shilov [375]. Shimura [347]. Short [86]. Siegel [212, 260]. Sierpinski [281, 390, 341, 297]. sieve [341]. sign [124].

Signed [256]. Similar [309]. Similarity [307, 330]. Simple [273, 145, 58]. Simplicial [400]. simplicity [373]. Simultaneous [60]. Singular [356, 120, 170]. Site [278]. small [366]. Smooth [277, 326, 366]. Sobolev [315]. Solutions [315, 181, 162, 56, 178, 121]. Some [186, 247, 331, 125, 321, 152, 18, 297, 202, 280, 80, 161, 162, 201, 378, 381, 117, 149]. Space [305, 259, 190, 101, 127, 34, 63, 221, 2, 132, 335, 153]. Space-Forms [305]. space-time [2]. Spaces [292, 253, 275, 276, 103, 373, 317, 107, 318, 179, 228, 185, 115, 194, 368, 227, 344, 8, 400, 353, 79, 336]. Spacetime [270]. Special [322, 327, 389]. Species [269]. Spectral [247, 264]. spectrum [120, 328]. sphere [141, 134, 356]. Spherical [45]. Spinor [127]. Square [286, 53]. squared [117]. squares [254, 117]. Squaring [53]. Stable [193, 283]. stably [330]. Stahl [363]. stamp [70]. Star [43]. Steenrod [254]. Steiner [332]. Steps [299]. Stiefel [289]. Störmer [77]. Strong [275, 388]. strongly [330]. structure [8]. study [70]. subadditive [173]. subalgebras [176]. Subdifferentials [263]. subdivisions [372]. Subgroups [15, 278]. Sublattices [309]. submerged [54]. Submodules [307]. Subquotients [302]. Subregular [303]. subsets [400]. subspaces [266, 335, 153]. Sum [320, 394, 82, 71]. Summability [195, 25, 109, 125]. Summation [283]. sums [366, 195, 88, 69]. supersingular [347]. supplementary [26]. Supporting [253]. supremum [174]. surface [374, 20]. surfaces [371, 49, 360, 99, 151, 191, 1]. Surjective [291]. symbols [392]. Symmetric [305, 266, 146, 119, 118, 317, 200, 211, 204, 166, 202, 13, 43]. symmetry [188, 4]. synthèse [132]. synthesis [132]. system [151]. systèmes [121]. Systems [307, 184, 315, 75, 332].

Tabling [239]. Tchebycheff [57]. Tensor [284, 287, 362, 189, 222, 196]. tensors [179]. ternary [144]. Tessellations [305]. Testing [278]. th [89]. their [332]. Theorem [270, 255, 212, 146, 123, 198, 75, 101, 96, 144, 114, 90, 32, 27, 42, 23]. théorème [96, 41]. theorems [175, 201, 25, 109, 125, 152, 57, 393]. theoretic [93]. Theory [294, 303, 278, 148, 49, 339, 327, 46, 113, 196, 41, 136, 106, 171, 76, 35]. third [217]. Three [65, 99, 30, 154, 19]. threefolds [376]. Thue [123]. Tietze [359]. timbres [70]. Time [239, 2]. Time-Tabling [239]. tool [49]. topological [103, 113]. Topos [246]. Tori [301]. toric [376]. Toroidal [274, 306]. torsion [362]. torsion-free [362]. torus [333]. Towers [264, 364]. Traces [291, 24]. Tractable [252]. Transcendental [389]. Transfer [283, 41]. Transferability [284]. transfert [41]. Transform [285, 375]. transformation [171]. transformations [224, 2, 48, 153]. transforms [155, 225, 147, 55, 97]. Translation [345]. treatment [351, 28, 4]. Trees [238, 240]. trigonometric [223]. trigonometrical [224]. triple [332]. twists [334]. Two [111, 371, 183, 33, 191, 342]. two-dimensional [183]. Type [270, 313, 285, 268, 224, 222, 185, 367, 51]. types [162, 188].

ultraspherical [197, 391]. Unified [76]. uniformly [344, 353]. Union [78]. Unipotent [278]. unique [348]. Uniqueness [269, 64, 16]. unit [146]. Unitary [302, 205, 134]. Univalent [271]. upon [3]. using [86].

V [96]. Valuations [372]. value [327, 189]. valued [190, 246]. Values [261, 389, 322, 353, 334, 369]. variables [69, 211, 188, 207]. variances [122]. variant [396]. Variation [221, 12, 68]. variational [326]. variations [374, 341, 156]. varieties [95, 338]. variety [347]. vector [371, 82, 8, 153]. vectorial [394]. Veränderlichen [49]. Verma [333]. vertex [111]. vertex-regular [111]. very [371]. via [335, 382]. viewpoint [128]. Viot [387]. Virasoro [274, 274]. Virasoro-Heisenberg [274]. Virasoro-Toroidal [274]. Vojta [386]. vortices [19]. Vorticity [92].

Waring [212]. Water [54]. Watson [55]. waves [54, 20]. weakly [317, 185]. weakly-compact [185]. Weighted [199, 275]. weights [336]. Weyl [26]. which [184, 167]. Whittaker [342]. whose [151]. Wilson [27]. Wold [287].

Zero [243, 358]. zero-divisors [358]. zeros [187, 257]. Zeta [134, 394, 26]. zetafunctions [394].

References

Lalan:1949:SCM

[1] Victor Lalan. Sur les surfaces à courbure moyenne isotherme. (French) [On surfaces with isotherm mean curvature]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):6–28, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schild:1949:DST

[2] Alfred Schild. Discrete space-time and integral Lorentz transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??): 29–47, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Turnbull:1949:NUG

[3] H. W. Turnbull. Note upon the generalized Cayleyan operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):48–56, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weyl:1949:EAT

[4] Hermann Weyl. Elementary algebraic treatment of the quantum mechanical symmetry problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):57–68, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fenchel:1949:CCF

[5] W. Fenchel. On conjugate convex functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):73–77, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahler:1949:CLA

[6] K. Mahler. On the critical lattices of arbitrary point sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):78–87, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruck:1949:NCF

[7] R. H. Bruck and H. J. Ryser. The nonexistence of certain finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):88–93, 1949. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Menger:1949:GVS

[8] Karl Menger. Generalized vector spaces. I. The structure of finitedimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):94–104, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1949:GRB

[9] Irving Kaplansky. Groups with representations of bounded degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):105–112, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1949:MCI

[10] H. W. Ellis. Mean-continuous integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):113–124, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snapper:1949:CIM

[11] Ernst Snapper. Completely indecomposable modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):125–152, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morse:1949:FBF

[12] Marston Morse and William Transue. Functionals of bounded Fréchet variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):153–165, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1949:DPI

[13] G. de B. Robinson. On the disjoint product of irreducible representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):166–175, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1949:BPF

[14] M. H. Stone. Boundedness properties in function-lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):176–186, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1949:SFI

[15] Marshall Hall, Jr. Subgroups of finite index in free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):187–190, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Titchmarsh:1949:UGF

[16] E. C. Titchmarsh. On the uniqueness of the Green’s function associated with a second-order differential equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):191–198, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Birkhoff:1949:NNF

[17] Garrett Birkhoff and Lindley Burton. Note on Newtonian force-fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):199– 208, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1949:SPE

[18] S. Minakshisundaram and Å. Pleijel. Some properties of the eigenfunctions of the Laplace-operator on Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):242–256, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Synge:1949:MTV

[19] J. L. Synge. On the motion of three vortices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):257–270, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weinstein:1949:SW

[20] Alexander Weinstein. On surface waves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):271–278, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Busemann:1949:AMI

[21] Herbert Busemann. Angular measure and integral curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):279–296, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1949:DRI

[22] S. D. Chowla and John Todd. The density of reducible integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):297–299, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taussky:1949:TLM

[23] Olga Taussky. On a theorem of Latimer and MacDuffee. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):300–302, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frame:1949:CRB

[24] J. S. Frame. Congruence relations between the traces of matrix powers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):303–304, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1949:DTM

[25] G. G. Lorentz. Direct theorems on methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):305–319, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1949:GEZ

[26] S. Minakshisundaram. A generalization of Epstein zeta functions. With a supplementary note by Hermann Weyl. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1 (??):320–327, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1949:ACF

[27] N. S. Mendelsohn. Applications of combinatorial formulae to generalizations of Wilson’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):328–336, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rado:1949:ATR

[28] R. Rado. Axiomatic treatment of rank in infinite sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):337–343, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pall:1949:RQF

[29] Gordon Pall. Representation by quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):344–364, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1949:GDT

[30] Robert Frucht. Graphs of degree three with a given abstract group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):365–378, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1949:FLE

[31] G. F. D. Duff. Factorization ladders and eigenfunctions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):379–396, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahler:1949:TLF

[32] K. Mahler. On a theorem of Liouville in fields of positive characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):397–400, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davenport:1949:PTP

[33] H. Davenport and G. Pólya. On the product of two power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):1–5, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacDuffee:1949:OMF

[34] C. C. MacDuffee. Orthogonal matrices in four-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):69–72, 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1949:MPG

[35] A. Einstein and L. Infeld. On the motion of particles in General Relativity Theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(1):209–241 + 1, 1949. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Seifert:1950:ONR

[36] Herbert Seifert and William Threlfall. Old and new results on knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):1–15, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1950:FOI

[37] H. B. Mann. On the field of origin of an ideal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):16–21, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toth:1950:EPR

[38] László Fejes Tóth. Extremum properties of the regular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):22–31, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levi:1950:FSC

[39] F. W. Levi. On frequencies and semicontinuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):32–43, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1950:FLF

[40] W. T. Tutte. The factorization of locally finite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):44–49, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lesieur:1950:TTA

[41] Léonce Lesieur. Un théorème de transfert d’un anneau abstrait à l’anneau des polynomes. (French) [A theory of the transfer of an abstract ring to a polynomial ring]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):50–65, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Selberg:1950:EPP

[42] Atle Selberg. An elementary proof of the prime-number theorem for arithmetic progressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):66–78, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Staal:1950:SDS

[43] R. A. Staal. Star diagrams and the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):79–92, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1950:CP

[44] S. Chowla and H. J. Ryser. Combinatorial problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):93–99, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prenowitz:1950:SGM

[45] Walter Prenowitz. Spherical geometries and multigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):100–119, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1950:BIG

[46] A. Einstein. The Bianchi identities in the generalized theory of gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):120–128, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1950:GHD

[47] P. A. M. Dirac. Generalized Hamiltonian dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):129–148, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snapper:1950:PLT

[48] Ernst Snapper. Periodic linear transformations of affine and projective geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):149–151, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behnke:1950:ERF

[49] Heinrich Behnke and Karl Stein. Elementarfunktionen auf Riemannschen Flächen als Hilfsmittel für die Funktionentheorie mehrerer Veränderlichen. (German) [Elementary functions on Riemann surfaces as a tool for the functional theory of multiple mutables]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):152–165, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zassenhaus:1950:EDA

[50] Hans Zassenhaus. An equation for the degrees of the absolutely irreducible representations of a group of finite order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):166–167, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanathan:1950:ICR

[51] K. G. Ramanathan. Identities and congruences of the Ramanujan type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):168–178, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1950:CSD

[52] C. A. B. Smith and W. T. Tutte. A class of self-dual maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):179–196, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1950:SS

[53] W. T. Tutte. Squaring the square. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):197–209, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heins:1950:WWC

[54] Albert E. Heins. Water waves over a channel of finite depth with a submerged plane barrier. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):210–222, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodspeed:1950:RBF

[55] F. M. Goodspeed. The relation between functions satisfying a certain integral equation and general Watson transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):223–237, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holzer:1950:MSD

[56] L. Holzer. Minimal solutions of Diophantine equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):238–244, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rademacher:1950:HTC

[57] Hans Rademacher and I. J. Schoenberg. Helly’s theorems on convex domains and Tchebycheff’s approximation problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):245–256, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jacobson:1950:EAS

[58] N. Jacobson. Enveloping algebras of semi-simple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):257–266, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1950:AAD

[59] D. H. Lehmer. Approximations to the area of an n-dimensional ellipsoid. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):267–282, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raisbeck:1950:SDA

[60] Gordon Raisbeck. Simultaneous Diophantine approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):283–288, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chatland:1950:EAR

[61] H. Chatland and H. Davenport. Euclid’s algorithm in real quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):289–296, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

James:1950:GII

[62] R. D. James. A generalized integral. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):297–306, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moor:1950:GDS

[63] Arthur Moór. Généralisation du scalaire de courbure et du scalaire principal d’un espace finslérien à n dimensions. (French) [Generalization of the scalar curvature and the principal scalar of a Finsler space in n dimensions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):307– 313, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sears:1950:NUG

[64] D. B. Sears. Note on the uniqueness of the Green’s functions associated with certain differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):314–325, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ApSimon:1950:TFR

[65] Hugh ApSimon. Three faciallyregular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):326–330 (2 plates), 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Todd:1950:NPR

[66] J. A. Todd. Note on a paper by Robinson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):331–333, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1950:IRI

[67] G. de B. Robinson. Induced representations and invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):334–343, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morse:1950:FVS

[68] Marston Morse and William Transue. The Fréchet variation, sector limits, and left decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):344–374, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kac:1950:DMP

[69] Mark Kac and Harry Pollard. The distribution of the maximum of partial sums of independent random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):375–384, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Touchard:1950:CED

[70] Jacques Touchard. Contribution à l’étude du problème des timbres poste. (French) [Contribution to the study of the postage-stamp problem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):385–398, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

LeVeque:1950:RSC

[71] W. J. LeVeque. On representations as a sum of consecutive integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):399–405, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Niven:1950:ICA

[72] Ivan Niven. The iteration of certain arithmetic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):406–408, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1950:IFO

[73] Rufus Isaacs. Iterates of fractional order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):409–416, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1950:LGA

[74] Robert Frucht. Lattices with a given abstract group of automorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):417– 419, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cramlet:1950:GTJ

[75] Clyde M. Cramlet. A generalization of a theorem of Jacobi on systems of linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):420–426, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1950:UFT

[76] Max Wyman. Unified field theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):427–439, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deVogelaere:1950:EHP

[77] René de Vogelaere. Équation de Hill et problème de Störmer. (French) [Hill’s equation and Störmer’s problem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):440–456, 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Springer:1950:UCH

[78] C. E. Springer. Union curves of a hypersurface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):457–460, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1950:IRM

[79] A. H. Stone. Incidence relations in multicoherent spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):461–480, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Valentine:1950:SPC

[80] F. A. Valentine. Some properties of C-convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):481–488, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:1950:QS

[81] J. W. Green and W. Gustin. Quasiconvex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):489–507, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:1950:VSC

[82] J. W. Green and W. Gustin. On the vector sum of continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):508–512, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1951:HFF

[83] P. A. M. Dirac. The Hamiltonian form of field dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):1–23, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Croisot:1951:ALD

[84] R. Croisot. Axiomatique des lattices distributives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):24–27, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sholander:1951:PDL

[85] Marlow Sholander. Postulates for distributive lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):28–30, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Byrne:1951:SFB

[86] Lee Byrne. Short formulations of Boolean algebra, using ring operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):31–33, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1951:ISG

[87] J. Lambek. The immersibility of a semigroup into a group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):34–43, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hua:1951:ESA

[88] Loo-Keng Hua. On exponential sums over an algebraic number field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 44–51, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porcelli:1951:PFS

[89] Pasquale Porcelli and Gordon Pall. A property of Farey sequences, with applications to qth power residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):52–53, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macbeath:1951:CTA

[90] A. M. Macbeath. A compactness theorem for affine equivalence-classes of convex regions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):54–61, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burchnall:1951:BP

[91] J. L. Burchnall. The Bessel polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):62–68, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Truesdell:1951:VA

[92] C. Truesdell. Vorticity averages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 69–86, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:ABAa

[93] David Ellis. Autometrized Boolean algebras. I. Fundamental distance-theoretic properties of B. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):87–93, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruck:1951:FNN

[94] R. H. Bruck. Finite nets. I. Numerical invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):94–107, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kodaira:1951:GFM

[95] Kunihiko Kodaira. Green’s forms and meromorphic functions on compact analytic varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):108–128, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dieudonne:1951:TLN

[96] Jean Dieudonné. Sur le théorème de Lebesgue–Nikodym. V. (French) [On the Lebesgue–Nikodym theorem. V]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):129– 139, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lukacs:1951:CFT

[97] Eugene Lukacs and Otto Szász. Certain Fourier transforms of distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):140–144, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:ABAb

[98] David Ellis. Autometrized Boolean algebras. II. The group of motions of B. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):145–147, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1951:RSG

[99] Patrick Du Val. On regular surfaces of genus three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):148–154, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hutcherson:1951:CIP

[100] W. R. Hutcherson. A cyclic involution of period eleven. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):155–158, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Derry:1951:DTC

[101] Douglas Derry. The duality theorem for curves of order n in n-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):159–163, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stanton:1951:MG

[102] R. G. Stanton. The Mathieu groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):164–174, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bing:1951:MTS

[103] R. H. Bing. Metrization of topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):175–186, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1951:DS

[104] L. M. Kelly. Distance sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):187–194, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Infeld:1951:RGE

[105] L. Infeld and A. E. Scheidegger. Radiation and gravitational equations of motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):195–207, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakayama:1951:NNG

[106] Tadasi Nakayama. Non-normal Galois theory for non-commutative and non-semisimple rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):208–218, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dowker:1951:CPS

[107] C. H. Dowker. On countably paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):219–224, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1951:R

[108] Robert Steinberg. The representations of GL(3, q), GL(4, q), PGL(3, q), and PGL(4, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):225–235, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1951:DTM

[109] G. G. Lorentz. Direct theorems on methods of summability. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):236–256, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heilbronn:1951:EAC

[110] H. Heilbronn. On Euclid’s algorithm in cyclic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):257–268, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ApSimon:1951:TVR

[111] Hugh ApSimon. Two vertex-regular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):269–271, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1951:CIP

[112] V. L. Klee, Jr. On certain intersection properties of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):272–275, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1951:TTF

[113] James A. Jenkins. On the topological theory of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):276–289, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1951:TDR

[114] Irving Kaplansky. A theorem on division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):290–292, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1951:DSO

[115] Israel Halperin and Hidegoro Nakano.

Discrete semi-ordered linear spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):293–298, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1951:GPC

[116] Gerald Berman. A generalization of the Pappus configuration. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):299–303, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Willcocks:1951:NSP

[117] T. H. Willcocks. A note on some perfect squared squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):304–308, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chung:1951:MRS

[118] J. H. Chung. Modular representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):309–327, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1951:RCS

[119] S. Chowla, I. N. Herstein, and W. K. Moore. On recursions connected with symmetric groups. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):328–334, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coddington:1951:NSS

[120] E. A. Coddington and N. Levinson. On the nature of the spectrum of singular second order linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):335–338, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reeb:1951:SPC

[121] Georges Reeb. Sur les solutions périodiques de certains systèmes différentiels perturbés. (French) [On the periodic solutions of certain perturbed differential equations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):339–362, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fraser:1951:NSL

[122] D. A. A. Fraser. Normal samples with linear constraints and given variances. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):363–366, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1951:TAT

[123] Alfred Brauer and R. L. Reynolds. On a theorem of Aubry–Thue. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):367–374, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1951:CSC

[124] Paul Erdős and Harold N. Shapiro.

On the changes of sign of a certain error function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):375–385, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macphail:1951:STA

[125] M. S. Macphail. Some theorems on absolute summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):386–390, 1951.

CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1951:EF

[126] H. S. M. Coxeter. Extreme forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):391–441, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hlavaty:1951:SSL

[127] Václav Hlavatý. Spinor space and line geometry. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):442–459, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bochner:1951:NVD

[128] S. Bochner. A new viewpoint in differential geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):460–470, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:DPA

[129] H. W. Ellis. Darboux properties and applications to non-absolutely convergent integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):471–485, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ankeny:1951:CNC

[130] N. C. Ankeny and S. Chowla. The class number of the cyclotomic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):486–494, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1951:CIM

[131] Marshall Hall and H. J. Ryser. Cyclic incidence matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):495–502, 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwartz:1951:ASH

[132] Laurent Schwartz. Analyse et synthèse harmoniques dans les espaces de distributions. (French) [Harmonic analysis and synthesis in distribution space]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):503–512, 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Touchard:1952:PCF

[133] Jacques Touchard. Sur un problème de configurations et sur les fractions continues. (French) [On a configuration problem and on continued fractions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):2–25, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1952:ZFU

[134] S. Minakshisundaram. Zeta functions on the unitary sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):26–30, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldhaber:1952:HMC

[135] J. K. Goldhaber. The homomorphic mapping of certain matric algebras onto rings of diagonal matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):31–42, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murdoch:1952:CNI

[136] D. C. Murdoch. Contributions to noncommutative ideal theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):43–57, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davenport:1952:NND

[137] H. Davenport and P. Erdős. Note on normal decimals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):58–63, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1952:PSG

[138] Henry B. Mann. On products of sets of group elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):64–66, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Simons:1952:FCM

[139] William H. Simons. The Fourier coefficients of the modular function λ(τ). Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):67–80, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gans:1952:AEG

[140] David Gans. Axioms for elliptic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):81–92, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feld:1952:GLE

[141] J. M. Feld. On the geometry of lineal elements on a sphere, Euclidean kinematics, and elliptic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):93–110, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothberger:1952:PPH

[142] Fritz Rothberger. On the property C and a problem of Hausdorff. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):111–116, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothberger:1952:RED

[143] Fritz Rothberger. A remark on the existence of a denumerable base for a family of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):117–119, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1952:EMT

[144] Burton W. Jones. An extension of Meyer’s theorem on indefinite ternary quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):120–128, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Albert:1952:SAR

[145] A. A. Albert. On simple alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):129–135, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beatty:1952:SPR

[146] S. Beatty and N. D. Lane. A symmetric proof of the Riemann–Roch theorem, and a new form of the unit theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):136–148, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1952:IT

[147] Charles Fox. Iterated transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):149– 161, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1952:TNC

[148] Christine Williams Ayoub. A theory of normal chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):162–188, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moor:1952:QRG

[149] Arthur Moór. Quelques remarques sur la généralisation du scalaire de courbure et du scalaire principal. (French) [Some remarks on the generalization of the scalar curvature and the principal scalar]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):189–197, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graves:1952:CCO

[150] Ross E. Graves. A closure criterion for orthogonal functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):198–203, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1952:SWC

[151] Patrick Du Val. On surfaces whose canonical system is hyperelliptic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):204– 221, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1952:STD

[152] Henry B. Mann. Some theorems on difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):222–226, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thrall:1952:GCB

[153] R. M. Thrall. On a Galois connection between algebras of linear transformations and lattices of subspaces of a vector space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):227–239, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1952:ORG

[154] Robert Frucht. A one-regular graph of degree three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):240–247, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1952:EFT

[155] G. F. D. Duff. F -equation Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):248–256, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karush:1952:IPC

[156] William Karush. Isoperimetric problems in the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):257–280, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karush:1952:IEA

[157] William Karush. The index of an extremal arc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):281–294, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1952:CAP

[158] A. J. Hoffman. Cyclic affine planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):295–301, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1952:FPG

[159] Gerald Berman. Finite projective geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):302–313, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1952:FG

[160] W. T. Tutte. The factors of graphs.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):314–328, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

White:1952:SCG

[161] Paul A. White. Some characterizations of generalized manifolds with boundaries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):329–342, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faircloth:1952:NSS

[162] Olin B. Faircloth. On the number of solutions of some general types of equations in a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):343–351, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanathan:1952:AQF

[163] K. G. Ramanathan. Abelian quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):352–368, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1952:GCC

[164] Franklin Haimo. Groups with a certain condition on conjugates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):369–372, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1952:CJH

[165] G. de B. Robinson. On a conjecture by J. H. Chung. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):373–380, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osima:1952:IRS

[166] Masaru Osima. On the irreducible representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):381–384, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevenson:1952:PFD

[167] A. F. Stevenson and W. A. Bassali. On the possible forms of differential equation which can be factorized by the Schrödinger–Infeld method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):385–395, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gelbaum:1952:MS

[168] B. R. Gelbaum and G. K. Kalisch. Measure in semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):396–406, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1952:CPF

[169] M. S. Robertson. A coefficient problem for functions regular in an annulus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):407–423, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glover:1952:AFI

[170] I. E. Glover. Analytic functions with an irregular linearly measurable set of singular points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):424–435, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1952:NRI

[171] P. G. Rooney. A new representation and inversion theory for the Laplace transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):436–444, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petersen:1952:NDS

[172] G. M. Petersen. A note on divergent series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):445–454, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1952:MSF

[173] G. G. Lorentz. Multiply subadditive functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):455–462, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1952:SFA

[174] Israel Halperin. The supremum of a family of additive functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):463–479, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1952:MCT

[175] G. A. Dirac. Map-colour theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):480–490, 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zemmer:1952:SFD

[176] Joseph L. Zemmer, Jr. On the subalgebras of finite division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):491–503, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duncan:1952:DLA

[177] D. G. Duncan. On D. E. Littlewood’s algebra of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):504–512, 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matte:1953:NSO

[178] Alphonse Matte. Sur de nouvelles solutions oscillatoires de équations de la gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):1–16, 1953. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:HTN

[179] G. F. D. Duff. Harmonic p-tensors on normal hyperbolic Riemannian spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):57–80, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1953:BAR

[180] R. S. Pierce. The Boolean algebra of regular open sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):95–100, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ewing:1953:CSF

[181] G. M. Ewing and W. R. Utz. Continuous solutions of the functional equation f^{n}(x) = f(x). Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):101–103, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rainville:1953:GFB

[182] E. D. Rainville. Generating functions for Bessel and related polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):104–106, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1953:TDB

[183] P. Butzer. On two-dimensional Bernstein polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):107–113, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschman:1953:SPD

[184] I. I. Hirschman, Jr. Systems of partial differential equations which generalize the heat equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):118–128, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grothendieck:1953:ALF

[185] A. Grothendieck. Sur les applications linéaires faiblement compactes d’espaces du type C(K). (French) [On the applications of linear weakly-compact spaces of the type C(K)]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):129–173, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bourgin:1953:SMF

[186] D. G. Bourgin. Some multiplicative functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):174–178, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boas:1953:IFN

[187] R. P. Boas, Jr. Integral functions with negative zeros. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):179–184, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Slepian:1953:NST

[188] David Slepian. On the number of symmetry types of Boolean functions of n variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):185–193, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:BVP

[189] G. F. D. Duff. Boundary value problems associated with the tensor Laplace equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):196–210, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arens:1953:ECP

[190] Richard Arens. Extension of coverings, of pseudometrics, and of linearspace-valued mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):211–215, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1953:RSG

[191] Patrick Du Val. Regular surfaces of genus two. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):216–237, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Churchhouse:1953:CI

[192] R. F. Churchhouse. A criterion for irrationality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):253–260, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1953:SL

[193] Harvey Cohn. Stable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):261–270, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1953:FS

[194] Israel Halperin. Function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):273– 288, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1953:SMD

[195] J. D. Hill. Summability methods defined by Riemann sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):289–296, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamson:1953:CTE

[196] K. W. Lamson. On the curvature tensor of Einstein’s generalized theory of gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):297–300, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brafman:1953:RBU

[197] Fred Brafman. A relation between ultraspherical and Jacobi polynomial sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):301–305, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1953:TG

[198] Leonard Carlitz. A theorem of Glaisher. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):306–316, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1953:WQP

[199] Leonard Carlitz. Weighted quadratic partitions over a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):317–323, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1953:ASD

[200] J. G. Elliott. Autometrization and the symmetric difference. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):324–331, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldhaber:1953:SMT

[201] J. K. Goldhaber and G. Whaples.

On some matrix theorems of Frobenius and McCoy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):332–335, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osima:1953:SRC

[202] Masaru Osima. Some remarks on the characters of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):336–343, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maranda:1953:AIR

[203] Jean-Marie Maranda. On B-adic integral representations of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):344–355, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagao:1953:NMR

[204] Hirosi Nagao. Note on the modular representations of symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):356–363, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1953:UGG

[205] G. C. Shephard. Unitary groups generated by reflections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):364–383, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1953:EDF

[206] H. S. M. Coxeter and J. A. Todd. An extreme duodenary form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):384–392, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tyler:1953:NIF

[207] G. W. Tyler. Numerical integration of functions of several variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):393–412, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1953:NEC

[208] S. S. Shrikhande. The non-existence of certain affine resolvable balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):413–420, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1953:CDS

[209] T. G. Ostrom. Concerning difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):421–424, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1953:RDS

[210] Emma Lehmer. On residue difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):425–432, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mordell:1953:NLS

[211] L. J. Mordell. Note on the linear symmetric congruence in n variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):433–438, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1953:WST

[212] R. G. Ayoub. On the Waring–Siegel theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):439–450, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watson:1953:IRP

[213] G. L. Watson. On integers n relatively prime to [αn]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):451–455, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Estermann:1953:NPL

[214] Theodor Estermann. On the number of primitive lattice points in a parallelogram. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):456–459, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sholander:1953:PBA

[215] Marlow Sholander. Postulates for Boolean algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):460–464, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Copeland:1953:CIB

[216] A. H. Copeland, Sr. and Frank Harary. A characterization of implicative Boolean rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):465–469, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldberg:1953:ELA

[217] S. I. Goldberg. Extensions of Lie algebras and the third cohomology group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):470–476, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1953:FSF

[218] D. G. Higman. Focal series in finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):477–497, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1953:FCG

[219] Franklin Haimo. The FC-chain of a group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):498–511, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1953:DPC

[220] N. D. Lane and Peter Scherk. Differentiable points in the conformal plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):512–518, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1953:VCC

[221] T. K. Pan. Variation of congruences of curves of an orthogonal ennuple in a Riemannian space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):519–523, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:TEE

[222] G. F. D. Duff. A tensor equation of elliptic type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):524–535, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walmsley:1953:NTS

[223] Charles Walmsley. Null trigonometric series in differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):536–543, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agarwal:1953:GTB

[224] R. P. Agarwal. General transformations of bilateral cognate trigonometrical series of ordinary hypergeometric type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):544–553, 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyson:1953:FTD

[225] F. J. Dyson. Fourier transforms of distribution functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):554–558, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1953:LCB

[226] P. L. Butzer. Linear combinations of Bernstein polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):559–567, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1953:RLF

[227] G. G. Lorentz and D. G. Wertheim. Representation of linear functionals on Köthe spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):568–575, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1953:FSD

[228] H. W. Ellis and Israel Halperin. Function spaces determined by a levelling length function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):576–592, 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agmon:1954:RML

[229] S. Agmon. The relaxation method for linear inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(3):382–392, 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Motzkin:1954:RML

[230] T. S. Motzkin and I. J. Schoenberg. The relaxation method for linear inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(3):393–404, 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1956:AE

[231] Leo Moser and Max Wyman. Asymptotic expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):225–233, 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ford:1956:MFT

[232] L. R. Ford and D. R. Fulkerson. Maximal flow through a network. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):399–404, 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1957:AEI

[233] Leo Moser and Max Wyman. Asymptotic expansions II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):194–209, 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1959:ABL

[234] Max Wyman. The asymptotic behavior of the Laurent coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):534–555, 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wigner:1964:PSN

[235] E. P. Wigner and M. M. Yanase. On the positive semidefinite nature of a certain matrix expression. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):397–406, 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmore:1964:CCG

[236] P. C. Gilmore and A. J. Hoffman. A characterization of comparability graphs and of interval graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 16(??):539–548, 1964. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1965:DIP

[237] L. Carlitz. The distribution of irreducible polynomials in several indeterminates. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):261–266, 1965. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edmonds:1965:PTF

[238] Jack Edmonds. Paths, trees, and flowers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 17(??):449–467, 1965. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dempster:1968:GCT

[239] M. A. H. Dempster. On the Gotlieb– Csima time-tabling algorithm. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(??):103– 119, 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Knuth:1968:AET

[240] Donald E. Knuth. Another enumeration of trees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 20(20):1077–1086, 1968. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). Check issue number??

Wyman:1970:AB

[241] M. Wyman and R. Wong. The asymptotic behaviour of μ(z, β, α). Canadian Journal of Mathematics = Journal canadien de mathématiques, 22 (??):1013–1023, 1970. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tromba:1976:ARM

[242] A. Tromba. Almost Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 28(??):640–652, 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geramita:1976:ODZ

[243] A. V. Geramita and J. H. Verner.

Orthogonal designs with zero diagonal. Canadian Journal of Mathematics = Journal canadien de mathématiques, XXVIII(1):215–224, 1976. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Avis:1980:ERM

[244] David Avis. On the extreme rays of the metric cone. Canadian Journal of Mathematics = Journal canadien de mathématiques, 32(1):126–144, 1980. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:1984:HMM

[245] J. M. Borwein and B. Richmond. How many matrices have roots? Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(2):286–299, 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higgs:1984:ITC

[246] D. Higgs. Injectivity in the topos of complete Heyting algebra valued sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(3):550–568, 1984. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cain:1984:SSP

[247] Bryan E. Cain. Some spectral properties of polar decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(6):973–985, 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nivat:1986:ERM

[248] Maurice Nivat and Dominique Perrin. Ensembles reconnaissables de mots biinfinis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(3):513–537, 1986. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bennett:1989:QIM

[249] F. Bennett. Quasigroup identities and Mendelsohn designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(??):341–368, 1989. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreck:1991:CMF

[250] M. Kreck and G. Triantafillou. On the classification of manifolds up to finite ambiguity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):356–370, 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allday:1999:CGK

[251] C. Allday and V. Puppe. On a conjecture of Goresky, Kottwitz and MacPherson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):3–9, 1999. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chacron:1999:TF

[252] M. Chacron, J.-P. Tignol, and A. R.

Wadsworth. Tractable fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):10–25, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabian:1999:SRS

[253] Marián Fabian and Boris S. Mordukhovich. Separable reduction and supporting properties of Fréchet-like normals in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):26–48, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ndombol:1999:AQC

[254] Bitjong Ndombol and M. El haouari.

Algèbres quasi-commutatives et carrés de Steenrod. (French) [Quasi-commutative algebras and Steenrod’s squares]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):

49–68, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reichstein:1999:THJ

[255] Zinovy Reichstein. On a theorem of Hermite and Joubert. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):69–95, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosler:1999:PCS

[256] Margit Rösler and Michael Voit. Partial characters and signed quotient hypergroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):96–116, 1999. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sauer:1999:MFP

[257] A. Sauer. Meromorphic functions with prescribed asymptotic behaviour, zeros and poles and applications in complex approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):117–129, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Savin:1999:DPA

[258] Gordan Savin and Wee Teck Gan. The dual pair G_{2} × PU_{3}(D) (p-adic case). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):130–146, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Suarez:1999:HAM

[259] Daniel Suárez. Homeomorphic analytic maps into the maximal ideal space of H^{∞}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):147–163, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tan:1999:PSE

[260] Victor Tan. Poles of Siegel Eisenstein series on U(n, n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):164–175, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderPoorten:1999:VDE

[261] Alfred van der Poorten and Kenneth S.

Williams. Values of the Dedekind eta function at quadratic irrationalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(1):176–224, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Betke:1999:AFL

[262] U. Betke and K. Böröczky, Jr. Asymptotic formulae for the lattice point enumerator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):225–249, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Combari:1999:CSC

[263] C. Combari, R. Poliquin, and L. Thibault.

Convergence of subdifferentials of convexly composite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):250–265, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deitmar:1999:SET

[264] Anton Deitmar and Werner Hoffman. Spectral estimates for towers of noncompact quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):266–293, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Enochs:1999:HQM

[265] Edgar E. Enochs and Ivo Herzog. A homotopy of quiver morphisms with applications to representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):294–308, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leung:1999:SSS

[266] Denny H. Leung and Wee-Kee Tang. Symmetric sequence subspaces of C(α), II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):309–325, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1999:ASO

[267] W. J. Martin and D. R. Stinson. Association schemes for ordered orthogonal arrays and (T, M, S)-nets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):326–346, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muhlherr:1999:EMQ

[268] Bernhard Mühlherr and Hendrik Van Maldeghem. Exceptional Moufang quadrangles of type F_{4}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):347–371, 1999.

CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mytnik:1999:UCS

[269] Leonid Mytnik. Uniqueness for a competing species model. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(2):372–448, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bahn:1999:BMT

[270] Hyoungsick Bahn and Paul Ehrlich. A Brunn–Minkowski type theorem on the Minkowski spacetime. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):449–469, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bshouty:1999:EUH

[271] D. Bshouty and W. Hengartner. Exterior univalent harmonic mappings with finite Blaschke dilatations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):470–487, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burgess:1999:HAS

[272] W. D. Burgess and Manuel Saorín. Homological aspects of semigroup gradings on rings and algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):488–505, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elduque:1999:PIE

[273] A. Elduque and A. V. Iltyakov. On polynomial invariants of exceptional simple algebraic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):506–522, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabbri:1999:RVH

[274] Marc A. Fabbri and Frank Okoh. Representations of virasoro-heisenberg algebras and Virasoro-toroidal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):523–545, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felten:1999:SCI

[275] M. Felten. Strong converse inequalities for averages in weighted L^{p} spaces on [−1, 1]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):546–565, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ferenczi:1999:QHI

[276] V. Ferenczi. Quotient hereditarily indecomposable Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):566–584, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mansfield:1999:SFD

[277] R. Mansfield, H. Movahedi-Lankarani, and R. Wells. Smooth finite dimensional embeddings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):585–615, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Panyushev:1999:PSA

[278] Dmitri I. Panyushev. Parabolic subgroups with Abelian unipotent radical as a testing site for invariant theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):616–635, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paul:1999:FOD

[279] Annegret Paul. First occurrence for the dual pairs (U(p, q), U(r, s)). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):636–657, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shumyatsky:1999:NSL

[280] Pavel Shumyatsky. Nilpotency of some Lie algebras associated with p-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):658–672, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barlow:1999:BMH

[281] Martin T. Barlow and Richard F. Bass.

Brownian motion and harmonic analysis on Sierpinski carpets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):673–744, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Echterhoff:1999:ICD

[282] Siegfried Echterhoff and John Quigg.

Induced coactions of discrete groups on C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):745–770, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Flicker:1999:SBP

[283] Yuval Z. Flicker. Stable bi-period summation formula and transfer factors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):771–791, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gratzer:1999:TPT

[284] G. Grätzer and F. Wehrung. Tensor products and transferability of semilattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):792–815, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1999:NFS

[285] Brian C. Hall. A new form of the Segal– Bargmann transform for Lie groups of compact type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):816–834, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:1999:LSM

[286] Henry H. Kim. Langlands–Shahidi method and poles of automorphic L-functions: Application to exterior square L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):835–849, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muhly:1999:TAI

[287] Paul S. Muhly and Baruch Solel. Tensor algebras, induced representations, and the Wold decomposition. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):850–880, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Witherspoon:1999:RRC

[288] Sarah J. Witherspoon. The representation ring and the centre of a Hopf algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(4):881–896, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Astey:1999:CCP

[289] L. Astey, S. Gitler, E. Micha, and G. Pastor. Cohomology of complex projective Stiefel manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):897–914, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balogh:1999:QCP

[290] Zoltán M. Balogh and Christoph Leuenberger. Quasiconformal contactomorphisms and polynomial hulls with convex fibers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):915–935, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

David:1999:GRN

[291] Chantal David, Hershy Kisilevsky, and Francesco Pappalardi. Galois representations with non-surjective traces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):936–951, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deitmar:1999:LMS

[292] Anton Deitmar and Werner Hoffmann. On limit multiplicities for spaces of automorphic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):952–976, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fisher:1999:EPN

[293] Stephen D. Fisher and Dmitry Khavinson. Extreme Pick–Nevanlinna interpolants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):977–995, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1999:CET

[294] Roger L. Jones, Joseph M. Rosenblatt, and Máté Wierdl. Counting in ergodic theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):996–1019, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kozlov:1999:FSM

[295] Dmitry N. Kozlov. On functions satisfying modular equations for infinitely many primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):1020–1034, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Litherland:1999:HAC

[296] R. A. Litherland. The homology of Abelian covers of knotted graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (5):1035–1072, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nielsen:1999:HPD

[297] Ole A. Nielsen. The Hausdorff and packing dimensions of some sets related to Sierpiński carpets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(5):1073–1088, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vakil:1999:CNQ

[298] Ravi Vakil. The characteristic numbers of quartic plane curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (5):1089–1120, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arnold:1999:FSL

[299] V. I. Arnold. First steps of local contact algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1123–1134, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arthur:1999:EFC

[300] James Arthur. Endoscopic L-functions and a combinatorial identity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1135–1148, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1999:LGG

[301] A. M. Cohen, H. Cuypers, and H. Sterk.

Linear groups generated by reflection tori. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1149–1174, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehrer:1999:RSU

[302] G. I. Lehrer and T. A. Springer. Reflection subquotients of unitary reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1175–1193, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lusztig:1999:SNE

[303] G. Lusztig. Subregular nilpotent elements and bases in K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1194–1225, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKay:1999:SAC

[304] John McKay. Semi-affine Coxeter– Dynkin graphs and G ⊆ SU_{2}(C). Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1226–1229, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartley:1999:STE

[305] Michael I. Hartley, Peter McMullen, and Egon Schulte. Symmetric tessellations on Euclidean space-forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1230–1239, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Monson:1999:RRT

[306] B. Monson and A. Ivić Weiss. Realizations of regular toroidal maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1240–1257, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baake:1999:SSR

[307] Michael Baake and Robert V. Moody.

Similarity submodules and root systems in four dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1258–1276, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1999:IIP

[308] G. C. Shephard. Isomorphism invariants for projective configurations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1277–1299, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conway:1999:ESS

[309] J. H. Conway, E. M. Rains, and N. J. A.

Sloane. On the existence of similar sublattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1300–1306, 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Johnson:1999:QIC

[310] Norman W. Johnson and Asia Ivić Weiss. Quadratic integers and Coxeter groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(6):1307–1336, 1999.

CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anonymous:1999:AII

[311] Anonymous. Author index — index des auteurs: for 1997 — pour 1997. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51 (6):1335–1339, 1999. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bauschke:2001:HPC

[312] H. H. Bauschke, O. Güler, A. S. Lewis, and H. S. Sendov. Hyperbolic polynomials and convex analysis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 53(3):470–488, 2001. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2004:FEA

[313] Jonathan M. Borwein, William F. Galway, and David Borwein. Finding and excluding b-ary Machin-type BBP formulae. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):1339–1342, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Anchouche:2010:ABC

[314] Boudjemâa Anchouche. On the asymptotic behavior of complete Kähler metrics of positive Ricci curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 3–18, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bouchekif:2010:SSE

[315] Mohammed Bouchekif and Yasmina Nasri. Solutions for semilinear elliptic systems with critical Sobolev exponent and Hardy potential. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):19–33, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Campbell:2010:BRR

[316] Peter S. Campbell and Monica Nevins.

Branching rules for ramified principal series representations of GL(3) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):34–51, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deng:2010:AAW

[317] Shaoqiang Deng. An algebraic approach to weakly symmetric Finsler spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (1):52–73, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ducrot:2010:PGE

[318] Arnaud Ducrot, Zhihua Liu, and Pierre Magal. Projectors on the generalized eigenspaces for neutral functional differential equations in L^{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 74–93, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:2010:LCG

[319] Wentang Kuo. The Langlands correspondence on the generic irreducible constituents of principal series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 94–108, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2010:SHM

[320] Chi-Kwong Li and Yiu-Tung Poon. Sum of Hermitian matrices with given eigenvalues: inertia, rank, and multiple eigenvalues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):109–132, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Makarov:2010:SAP

[321] Konstantin A. Makarov and Anna Skripka. Some applications of the perturbation determinant in finite von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):133–156, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Masri:2010:SVC

[322] Riad Masri. Special values of class group L-functions for CM fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 157–181, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prajs:2010:MAD

[323] Janusz R. Prajs. Mutually aposyndetic decomposition of homogeneous continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):182–201, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tang:2010:IEP

[324] Lin Tang. Interior h^{1} estimates for parabolic equations with LMO coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):202–217, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xing:2010:GDC

[325] Yang Xing. The general definition of the complex Monge–Ampère operator on compact Kähler manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 218–239, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Azagra:2010:SOS

[326] Daniel Azagra and Robb Fry. A second order smooth variational principle on Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):241–260, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chiang:2010:EVD

[327] Yik-Man Chiang and Mourad E. H. Ismail. Erratum to: On value distribution theory of second order periodic ODEs, special functions and orthogonal polynomials [Cno. 2245272]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):261, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goresky:2010:SEC

[328] Mark Goresky and Robert MacPherson. On the spectrum of the equivariant cohomology ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):262–283, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grbic:2010:SML

[329] Jelena Grbić and Stephen Theriault. Self-maps of low rank Lie groups at odd primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):284–304, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

He:2010:ASC

[330] Hua He, Yunbai Dong, and Xianzhou Guo. Approximation and similarity classification of stably finitely strongly irreducible decomposable operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 305–319, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jerrard:2010:SRR

[331] Robert L. Jerrard. Some rigidity results related to Monge–Ampère functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):320–354, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kral:2010:CRS

[332] Daniel Král, Edita Máčajová, Attila Pór, and Jean-Sébastien Sereni. Characterisation results for Steiner triple systems and their application to edgecolourings of cubic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 355–381, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lu:2010:VMQ

[333] Rencai Lü and Kaiming Zhao. Verma modules over quantum torus Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):382–399, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prasanna:2010:APC

[334] Kartik Prasanna. On p-adic properties of central L-values of quadratic twists of an elliptic curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):400–414, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sun:2010:CRS

[335] Shunhua Sun, Dechao Zheng, and Changyong Zhong. Classification of reducing subspaces of a class of multiplication operators on the Bergman space via the Hardy space of the bidisk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 415–438, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sundhall:2010:HFH

[336] Marcus Sundhäll and Edgar Tchoundja.

On Hankel forms of higher weights: the case of Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):439–455, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:2010:CSF

[337] Tonghai Yang. The Chowla–Selberg formula and the Colmez conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 456–472, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yun:2010:GMC

[338] Zhiwei Yun. Goresky–MacPherson calculus for the affine flag varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 473–480, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casals-Ruiz:2010:EAG

[339] Montserrat Casals-Ruiz and Ilya V. Kazachkov. Elements of algebraic geometry and the positive theory of partially commutative groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):481–519, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eriksen:2010:CND

[340] Eivind Eriksen. Computing noncommutative deformations of presheaves and sheaves of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):520–542, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hare:2010:MVS

[341] Kevin G. Hare. More variations on the Sierpiński sieve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):543–562, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishii:2010:WFR

[342] Taku Ishii. Whittaker functions on real semisimple Lie groups of rank two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):563–581, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Konyagin:2010:DP

[343] Sergei V. Konyagin, Carl Pomerance, and Igor E. Shparlinski. On the distribution of pseudopowers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):582–594, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martinez:2010:LUR

[344] J. F. Martínez, A. Moltó, J. Orihuela, and S. Troyanski. On locally uniformly rotund renormings in C(K) spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):595–613, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pronk:2010:TGO

[345] Dorette Pronk and Laura Scull. Translation groupoids and orbifold cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):614–645, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rupp:2010:R

[346] R. Rupp and A. Sasane. Reducibility in A_{R}(K), C_{R}(K), and A(K). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3): 646–667, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vollaard:2010:SLS

[347] Inken Vollaard. The supersingular locus of the Shimura variety for GU(1, s). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):668–720, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boocher:2010:FFU

[348] Adam Boocher, Michael Daub, Ryan K. Johnson, H. Lindo, S. Loepp, and Paul A. Woodard. Formal fibers of unique factorization domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 721–736, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ditzian:2010:ADA

[349] Z. Ditzian and A. Prymak. Approximation by dilated averages and K-functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):737–757, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dolinar:2010:GPQ

[350] Gregor Dolinar and Bojan Kuzma. General preservers of quasi-commutativity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (4):758–786, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Landquist:2010:ETC

[351] E. Landquist, P. Rozenhart, R. Scheidler, J. Webster, and Q. Wu. An explicit treatment of cubic function fields with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):787–807, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Legendre:2010:ELE

[352] Eveline Legendre. Extrema of low eigenvalues of the Dirichlet–Neumann Laplacian on a disk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):808–826, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ouyang:2010:BFC

[353] Caiheng Ouyang and Quanhua Xu.

BMO functions and Carleson measures with values in uniformly convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (4):827–844, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Samei:2010:BPA

[354] Ebrahim Samei, Nico Spronk, and Ross Stokke. Biflatness and pseudoamenability of Segal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 845–869, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Valdimarsson:2010:BLP

[355] Stefán Ingi Valdimarsson. The Brascamp–Lieb polyhedron. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 870–888, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xia:2010:SIO

[356] Jingbo Xia. Singular integral operators and essential commutativity on the sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):889–913, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zorn:2010:RPS

[357] Christian Zorn. Reducibility of the principal series for S̃p_{2}(F) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):914–960, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aleman:2010:MII

[358] Alexandru Aleman, Peter Duren, María J. Martín, and Dragan Vukotić. Multiplicative isometries and isometric zero-divisors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):961–974, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bjorndahl:2010:RTN

[359] Christina Bjorndahl and Yael Karshon.

Revisiting Tietze–Nakajima: local and global convexity for maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 975–993, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breslin:2010:CBS

[360] William Breslin. Curvature bounds for surfaces in hyperbolic 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (5):994–1010, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buckingham:2010:FCF

[361] Paul Buckingham and Victor Snaith.

Functoriality of the canonical fractional Galois ideal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1011–1036, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calvino-Louzao:2010:RET

[362] E. Calviño-Louzao, E. García-Río, and R. Vázquez-Lorenzo. Riemann extensions of torsion-free connections with degenerate Ricci tensor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1037– 1057, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2010:CS

[363] Yichao Chen and Yanpei Liu. On a conjecture of S. Stahl. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1058–1059, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Darmon:2010:HPT

[364] Henri Darmon and Ye Tian. Heegner points over towers of Kummer extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1060–1081, October 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Godinho:2010:FGM

[365] Leonor Godinho and M. E. Sousa-Dias. The fundamental group of S^{1}-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (5):1082–1098, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldmakher:2010:CSS

[366] Leo Goldmakher. Character sums to smooth moduli are small. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1099–1115, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jin:2010:DLO

[367] Yongyang Jin and Genkai Zhang. Degenerate p-Laplacian operators and Hardy type inequalities on H-type groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1116–1130, October 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleppe:2010:MSR

[368] Jan O. Kleppe. Moduli spaces of reflexive sheaves of rank 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1131–1154, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Young:2010:MCV

[369] Matthew P. Young. Moments of the critical values of families of elliptic curves, with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1155–1181, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yue:2010:FFR

[370] Hong Yue. A fractal function related to the John–Nirenberg inequality for Q_{α}(R^{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1182–1200, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alzati:2010:CVA

[371] Alberto Alzati and Gian Mario Besana. Criteria for very ampleness of rank two vector bundles over ruled surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1201–1227, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ardila:2010:VMP

[372] Federico Ardila, Alex Fink, and Felipe Rincón. Valuations for matroid polytope subdivisions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1228– 1245, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chaput:2010:QCM

[373] P. E. Chaput, L. Manivel, and N. Perrin. Quantum cohomology of minuscule homogeneous spaces III. Semisimplicity and consequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1246– 1263, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2010:HVM

[374] Jingyi Chen and Ailana Fraser. Holomorphic variations of minimal disks with boundary on a Lagrangian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1264–1275, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wassouli:2010:GPT

[375] Fouzia El Wassouli. A generalized Poisson transform of an L^{p}-function over the Shilov boundary of the ndimensional Lie ball. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1276– 1292, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kasprzyk:2010:CTF

[376] Alexander M. Kasprzyk. Canonical toric Fano threefolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1293– 1309, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2010:IHA

[377] Kyu-Hwan Lee. Iwahori–Hecke algebras of SL_{2} over 2-dimensional local fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1310–1324, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mo:2010:SEC

[378] Xiaohuan Mo and Changtao Yu. On some explicit constructions of Finsler metrics with scalar flag curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1325– 1339, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moeglin:2010:HOE

[379] C. Mœglin. Holomorphie des opérateurs d’entrelacement normalisés à l’aide des paramètres d’Arthur. (French) [Holomorphism of normalized interlacing operators with the help of Arthur parameters]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1340–1386, December 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pamuk:2010:HSE

[380] Mehmetcik Pamuk. Homotopy selfequivalences of 4-manifolds with free fundamental group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1387– 1403, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saroglou:2010:CES

[381] Christos Saroglou. Characterizations of extremals for some functionals on convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1404–1418, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:2010:BEM

[382] Dachun Yang and Dongyong Yang. BMO-estimates for maximal operators via approximations of the identity with non-doubling measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1419– 1434, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banica:2011:FBL

[383] T. Banica, S. T. Belinschi, M. Capitaine, and B. Collins. Free Bessel laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 3–37, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brudern:2011:AFP

[384] Jörg Brüdern and Trevor D. Wooley. Asymptotic formulae for pairs of diagonal cubic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):38–54, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chau:2011:PRF

[385] Albert Chau, Luen-Fai Tam, and Chengjie Yu. Pseudolocality for the Ricci flow and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 55–85, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2011:VC

[386] Xi Chen. On Vojta’s 1 + ϵ conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):86–103, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feng:2011:RIF

[387] Shui Feng, Byron Schmuland, Jean Vaillancourt, and Xiaowen Zhou. Reversibility of interacting Fleming–Viot processes with mutation, selection, and recombination. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):104–122, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Granirer:2011:SES

[388] Edmond E. Granirer. Strong and extremely strong Ditkin sets for the Banach algebras A_{p}^{r}(G) = A_{p} ∩ L^{r}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (1):123–135, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gun:2011:TNS

[389] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Transcendental nature of special values of L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 136–152, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hambly:2011:AFA

[390] B. M. Hambly. Asymptotics for functions associated with heat flow on the Sierpinski carpet. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):153–180, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:2011:CCD

[391] Mourad E. H. Ismail and Josef Obermaier. Characterizations of continuous and discrete q-ultraspherical polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):181–199, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:2011:EPE

[392] Mizan Rahman. An explicit polynomial expression for a q-analogue of the 9-j symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):200–221, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:2011:LTA

[393] Jiun-Chau Wang. Limit theorems for additive conditionally free convolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (1):222–240, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Essouabri:2011:MZF

[394] Driss Essouabri, Kohji Matsumoto, and Hirofumi Tsumura. Multiple zetafunctions associated with linear recurrence sequences and the vectorial sum formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):241–276, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ghate:2011:LIG

[395] Eknath Ghate and Vinayak Vatsal. Locally indecomposable Galois representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):277–297, April 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gun:2011:VLC

[396] Sanoli Gun and V. Kumar Murty.

A variant of Lehmer’s conjecture, II: the CM-case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):298–326, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jantzen:2011:DSA

[397] Chris Jantzen. Discrete series for padic SO(2n) and restrictions of representations of O(2n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):327–380, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ji:2011:CCA

[398] Kui Ji and Chunlan Jiang. A complete classification of AI algebras with the ideal property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):381–412, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Konvalinka:2011:GFH

[399] Matjaž Konvalinka and Mark Skandera. Generating functions for Hecke algebra characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):413–435, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mine:2011:SCO

[400] Kotaro Mine and Katsuro Sakai. Simplicial complexes and open subsets of nonseparable LF-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):436–459, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pavlicek:2011:MCM

[401] Libor Pavlíček. Monotonically controlled mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):460–480, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).